Growth Tech Inc., has earned $ 40 as Earnings Per Share. It proposes to pay $10 as dividend and reinvest the remaining. The return on investment by the firm is 20%. What is the growth rate of the firm’s earnings and dividends assuming a constant payout ratio and return on investment?


1
Expert's answer
2021-05-07T14:20:51-0400

we will find it by the formula:

"g=ROE\times(1-DPR)=20(1-\frac{10}{40})=15"